# ShoppingComp: Are LLMs Really Ready for Your Shopping Cart?

### v1

@@ C.1 Model Answer Prompt
## Output Format
Use XML format for the output.
<answer> <content>Analysis process</content> <candidate_product_list> <!-- Candidate product information --> <product> <product_name>Candidate product name</product_name> <check_list> <check_item> <demand>An atomic requirement broken down from the user’s needs</demand> <reason>Analysis of whether the product meets this atomic requirement</reason> <is_satisfied>Whether the product meets this atomic requirement, value is Yes or No</is_satisfied> </check_item> ... </check_list> </product> <product> ... </product> </candidate_product_list> <best>All individual items that meet the user’s needs, separated by commas. If no suitable product is found, this should be empty.</best> </answer>
Note: The product names provided must be complete, not partial names.
@@
C.2 report verify Prompt
C.2.1 SoP Judge Prompt
The following prompt is used to compute the Satisfaction of Products (SoP) metric.
@@ C.2.1 SoP Judge Prompt
You are an expert evaluator for question,rubric and product matching in an e-commerce. The "question" describes the user’s needs, the "rubric" provides the evaluation criteria, and the "product" is a single candidate item. Your task is to determine whether this product meets the user’s needs according to the rubric.
@@
Your final judgment must follow this format:
<answer>Satisfied / Not Satisfied /Unable to Determine</answer>
@@
1. Evaluate the product only as sold; do not combine products or add accessories.
@@
Agent-generated demands:
\${demands}
C.2.2 Scenario Coverage Judge Prompt
@@
You are a professional shopping analyst. You are given a real user shopping question and its corresponding set of rubrics. The agent has decomposed the question into a list of user demands. Your task is to determine whether the agent’s demands are all correct and fully cover every required point. The rubric set is provided for your reference.
## Evaluation criteria:
Each demand must be semantically accurate and collectively cover all requirements in the rubrics set.
## Output Format:
<analysis> Provide a detailed analysis of how each agent demand aligns with the rubric items, explaining which parts are correct, missing, or incorrect. </analysis>
<result> If all demands are correct and none are missing, output "correct"; If all demands are correct but some are missing, output "missing"; If some demands are incorrect but none are missing, output "incorrect"; If some demands are both incorrect and missing, output "fail". </result> USER:
@@
Agent-generated demands:
\${demands}
C.2.3 RV Judge Prompt
@@
You must evaluate the Reason according to the following criteria: 1. Whether the Reason correctly states the Product’s attributes, specifications, features, or functions.
@@
Your response must follow this exact format
@@
<result>1</result> <!-- 1 means the reasoning is fully accurate; 0 means factual errors are present -->
@@
C.2.4 Safety-Critical Judge Prompt
@@
<answer>correct</answer> or <answer>incorrect</answer> Provide a brief explanation after the <answer></answer> tag, clarifying the reason for your judgment.
@@
Appendix D Reliability Ablations for LLM-as-a-Judge
Model Gemini-2.5-Pro judge GPT-4o judge Deepseek-v3.1 judge
mean std mean std mean std
deepseek-v3.1 thinking 37.00% 1.97% 35.33% 2.72% 30.59% 2.74%
gemini-2.5-pro 47.79% 3.94% 43.44% 3.33% 36.13% 3.43%
gpt-5 50.13% 3.42% 46.23% 4.06% 38.55% 2.30%
Table 5: Soft accuracy (Selection Accuracy SoP) of different models under three automatic judges.
Prior work has shown that LLM evaluators often assign higher ratings to outputs generated by the same model family, a phenomenon known as self-preference bias  Panickssery et al. (2024). In our study, cross-judge comparison suggests that the Gemini-2.5-Pro judge exhibits minimal such bias: under the Gemini-judge setting, Gemini-2.5-Pro’s own score (47.79%) is lower than GPT-5’s (50.13%), and the same trend holds across other judges (GPT-4o: 46.23% vs. 43.44%; Deepseek: 38.55% vs. 36.13%). The absence of systematic self-favoring across independent judges indicates that Gemini-2.5-Pro does not inflate its own evaluations.
We further examined human–LLM agreement and found Gemini-2.5-Pro to be the most consistent judge, achieving the highest alignment rate (0.756%), outperforming GPT–4o (73%) and DeepSeek-v3.1 (68%). This observation aligns with RESEARCH RUBRICS  Sharma et al. (2025), which also reports Gemini-2.5-Pro as the most human-aligned and reliable automatic evaluator.
Appendix E Acknowledgments
We sincerely thank all contributors for their valuable collaboration and support throughout the development of ShoppingComp, including product operations, domain experts, and annotators.
E.1 Product and Operations
Huang Hui, Li Yuemeng, Li Muzhi, Lv Ningtao, Liang Sinian, Chen Jingzhe, Dong Chao, Chen Xuyang (annotation platform support), Peng Yuting (annotation platform support)
E.2 Shopping Experts
Mai Juanshu, Yang Jiahao, Wang Zhaonan, Qiao Yi, Lin Guansheng, Zhang Cheng, Ai Yuankun, Qian Jiawei, Wang Jinyu, Li Min, Zhou Qunwei, Guo Qian
E.3 Annotators
Lai Qinglan, Wang Bowen, Lu Lu, Qu Yaping, Chen Minmin, Cao Tian, Dai Xinping, Zhang Ming, Song Yanling, Lu Lin, Su Huiyu, Zhang Linyu, Lin Jie, Mou Honglei, Hu Huaijuan, Li Minghui
We also extend our sincere appreciation to the anonymous shopping experts for their domain knowledge, constructive feedback, and meticulous evaluation efforts.
Their contributions were essential to the success of this work, and we deeply appreciate their professionalism and dedication.

### v2

@@ F.1 Model Answer Prompt
- The product names provided must be complete, not partial names.
@@
## Output Format Use XML format for the output.
<answer>
  <content>Analysis process</content>
  <candidate_product_list>
    <!-- Candidate product information -->
    <product>
      <product_name>Candidate product name</product_name>
      <check_list>
        <check_item>
          <demand>An atomic requirement broken down from the user’s
          needs </demand>
          <reason>Analysis of whether the product meets this atomic
          equirement </reason>
          <is_satisfied>
          Whether the product meets this atomic requirement,
          value is Yes or No
          </is_satisfied>
        </check_item>
      </check_list>
    </product>
    <product>
    </product>
  </candidate_product_list>
  <best>
  All individual items that meet the user’s needs,
  separated by commas. If no suitable product is found,
  this should be empty.
  </best>
</answer>
@@
F.2 report verify Prompt
F.2.1 SoP Judge Prompt
The following prompt is used to compute the Score of Products (SoP) metric.
@@ F.2.1 SoP Judge Prompt
You are an expert evaluator for question,rubric and product matching in e-commerce.
The "question" describes the user’s needs, the "rubric" provides the evaluation criteria, and the "product" is a single candidate item.
Your task is to determine whether this product meets the user’s needs according to the rubric.
@@
1. Do not combine products or add accessories.
@@
## Output Format
<answer>Satisfied / Not Satisfied /Unable to Determine</answer>
<evidence>Explanation of the reasoning process with supporting source URLs.</evidence>
@@
F.2.2 Scenario Coverage Judge Prompt
@@
You are a **Strict Semantic Consistency Evaluator for Product Scenarios**. You are given an original real user question (‘question‘), a set of manually decomposed real user requirement scenarios (‘scene_set‘), and a single user requirement (‘demand‘) produced by a large language model. Your task is to determine whether the model-generated ‘demand‘ matches the original real user requirements in ‘scene_set‘.
## Judgment Criteria:
You are a **Strict Semantic Consistency Evaluator for Product Scenarios**. You are given an original real user question (‘question‘), a set of manually decomposed real user requirement scenarios (‘scene_set‘), and a single user requirement (‘demand‘) produced by a large language model. Your task is to determine whether the model-generated ‘demand‘ matches the original real user requirements in ‘scene_set‘. 1. For the requirement expressed in ‘demand‘, determine whether it is explicitly, unambiguously, and without any requirement expansion or reduction supported by at least one scenario in ‘scene_set‘.
2. The demand may perform reasonable and equivalent inference over the question and scene (e.g., inferring necessary product functions or parameters), but must not perform scenario scaling on the original problem. Any scaling that makes the required product functionality broader or narrower, harder or easier, is not allowed. For example, if the original problem is “horseback riding” and the inferred demand is “riding”, it should be judged as Not Matched.
3. If a single ‘demand‘ contains multiple constraints, all constraints must be fully covered by the same scenario(s) to be considered a match.
Carefully reason through this: do not introduce any requirements that did not appear in the original user question, especially if they would increase the difficulty of purchasing a suitable product.
4. Vague or placeholder-type requirements (e.g., "other requirements") must always be judged as not matched.
5. A scenario in ‘scene_set‘ must be clearly and concretely described in ‘demand‘, rather than being loosely summarized (e.g., generic phrases like "easy to use" or "stylish design" are insufficient), as such summarization may lead to purchasing a product that fails to meet the user’s actual detailed needs.
6. If the meaning of ‘demand‘ is that the product was released before September 2025, it should be considered matched.
## Output Format
<analysis> Provide a detailed analysis of the relationship between the user requirement and each scenario, explaining why it matches or does not match. </analysis>
<result> If it matches, output 1; if it does not match, output 0 </result>
USER:
@@
Model-Provided User Requirement (demand):
\${demand}
SYSTEM:
You are a “strict evaluator for product scene–demand consistency”. I have a real user question. From this question, I extracted one genuine user requirement scene, and I also asked a large language model to decompose the user needs into a set of demands (demand_set). Your task is to determine whether the extracted real requirement scene is covered by the demand_set provided by the model.
## Judgment Criteria:
1. The semantic meaning of the scene must be a subset of the semantic meaning of the demand_set.
2. All core functional and environmental constraints in the scene must have corresponding equivalent expressions in the demand_set. If any core constraint in the scene is missing in the demand_set, the scene must be judged as **Not Matched**. 3. One scene may correspond to multiple demands.
4. Vague or placeholder demands (e.g., “other requirements”) must always be judged as Not Matched.
5. When the scene specifies functional requirements for the product, the demand_set must contain corresponding functional requirements. The demands must not be overly strict or overly loose compared to the scene; otherwise, the result is Not Matched.
## Output Format
<analysis> Provide a detailed explanation of why the scene is matched or not matched. </analysis>
<result> If it matches, output 1; if it does not match, output 0 </result>
USER:
Question:
\${question}
Reference rubric:
\${rubric}
Model-Provided User Requirements (demand_set):
\${demand_set}
F.2.3 RV Judge Prompt
@@
You must evaluate the Reason according to the following criteria:
1. Whether the Reason correctly states the Product’s attributes, specifications, features, or functions.
@@
5. Whether the reasoning is logically sound, coherent, and free of logical flaws.
6. Whether the user can infer that the product satisfies their demand based on factual product information.
@@
<result>1</result>
<!-- 1 means the reasoning is fully accurate; 0 means factual errors are present -->
@@
F.2.4 Safety-Critical Judge Prompt
@@
<answer>correct</answer> or <answer>incorrect</answer>
<reasonCclarifying the reason for your judgment. </reason>